OpenStudy (anonymous):

A number is divided by 3. If 8 is subtracted from the quotient you get -6.

OpenStudy (anonymous):

\[\frac{n}{3}-8=6\] is what they want you to solve

OpenStudy (timmyk14):

6/3 - 8 = -6 2 - 8 = -6 -6 = -6

OpenStudy (anonymous):

A number is multiplied by 6. If 8 is subtracted from the product you get 34.

OpenStudy (anonymous):

Help me with these please, I'm not that great at word problems

OpenStudy (anonymous):

Answers

OpenStudy (anonymous):

These just baffle me.

OpenStudy (timmyk14):

6x - 8 = 34 6x = 42 x = 7

OpenStudy (anonymous):

But if you would to teach me how that is cool too :)

OpenStudy (anonymous):

A number is multiplied by 3. If 3 is subtracted from the product you get 24.

OpenStudy (timmyk14):

LOL, I'm dominating. LOVE MATH!!

OpenStudy (anonymous):

A number is divided by 2. If 16 is added to the quotient you get 20.

OpenStudy (anonymous):

A number is multiplied by 5. If 12 is added to the product you get 37.

OpenStudy (anonymous):

Inequalitys?

OpenStudy (timmyk14):

x/2 + 16 = 20 x/2 = 4 8 = 4

OpenStudy (anonymous):

-3x - 2 > 7

OpenStudy (timmyk14):

-3x - 2 > 7 -3x > 9 x < -3

OpenStudy (anonymous):

7x + 2 > 30

OpenStudy (timmyk14):

7x + 2 > 30 7x > 28 x = 4

OpenStudy (anonymous):

-9x - 3 > 51

OpenStudy (timmyk14):

-9x - 3 > 51 -9x > 54 x < -6

OpenStudy (anonymous):

8x + 6 54

OpenStudy (timmyk14):

What's the inequality symbol?

OpenStudy (anonymous):

-5x + 7 > 42

OpenStudy (timmyk14):

-5x + 7 > 42 -5x > 35 x < -7

OpenStudy (anonymous):

-3x - 2 > 7

OpenStudy (timmyk14):

-3x - 2 > 7 -3x > 9 x < -3

OpenStudy (timmyk14):

Can I have a medal please.

OpenStudy (anonymous):

Hold on i will have more for you in a sec please if you still want to help

OpenStudy (timmyk14):

Okay :) These are the easy ones ;) I just learned about inequalities maybe two days ago

OpenStudy (anonymous):

8 − t = 6

OpenStudy (timmyk14):

What does that say

OpenStudy (anonymous):

8 - t = 6

OpenStudy (anonymous):

11y = 88

OpenStudy (timmyk14):

8

OpenStudy (timmyk14):

That all @Tk12345432 ?

OpenStudy (anonymous):

21 + p = 42

OpenStudy (anonymous):

What is the slope of the line that contains the points (-3, -2) and (4, -2)

OpenStudy (timmyk14):

42 - 21 = p 21 = 21 p = 21

OpenStudy (anonymous):

Can you do slopes

OpenStudy (timmyk14):

It has no slope. It's a straight line.

OpenStudy (timmyk14):

Wait..(x,y). X is the horizontal line, and Y is the vertical line. Sorry I was wrong...

OpenStudy (anonymous):

Im still lost

OpenStudy (timmyk14):

Wait, I'm confused.. -3 is on the x-axis. So is 4. The y plots for these coordinates are the same. So basically there is no slope. Look at a drawing..

OpenStudy (timmyk14):

 Okay, I'm also a bit lost..Are these two seperate lines? Or is this supposed to make one line?